A straight rod of length L = 24.8 cm, carries a uniform charge, ? = 2.55
Answer #1

dq = λ dx x . Charge per unit length: λ = Q/L Charge on slice da: dq Adz dE Electric field generated by slice da: dE Electric
Answer #2
Ey = kQ/( d * sqrt(L² + d²) ).

Ey = (kQ/L) * ? d/( (d² + x²)³/2) dx, with the limits of integration being 0 to L. That little exponent is 3/2, in case it's difficult to read. Please note that the Q/L term is the linear charge density.

Since I saw the form d² + x², I used trigonometric substitution to make x = dtan(u) and dx = sec²(u).
Answer #3
DELETED